C=0.01Q^3+0.5Q^2+Q+1000

derive

Fixed cost

variable cost

average fixed cost

average variable cost


1
Expert's answer
2021-10-06T12:04:31-0400

"C=0.01Q^3+0.5Q^2+Q+1000"average variable cost


a) Fixed cost.

This refers to the cost which does not change when there is either an increase or decrease in the number of goods or services sold or produced.


Hence,

"Fixed\space Cost= 1000"


b) Variable cost.

This refers to the cost of all variable inputs in a production process.


Hence,

"Variable\space Cost\space (VC)\space function= 0.01Q^3+0.5Q^2+Q"


c) Average fixed cost.

This is calculated by dividing fixed cost (FC) by the quantity (Q).


Hence,

"AFC=\frac {FC}{q}"


"Average\space fixed\space cost\space (AFC)= \frac {1000}{q}"


d) Average variable cost.

This is the firm's variable cost per unit of output. In many cases, the function will start decreasing and then increase at one point.


Hence,

"Average\space variable\space cost =\frac {VC}{q}"


"=\frac{0.01Q^3+0.5Q^2+Q}{Q}"


"=0.01Q^2+0.5Q+1"